THE 4^{3}Π_{g} STATE OF Na_{2}: VIBRATIONAL NUMBERING AND HYPERFINE STRUCTURE

\underline{GUENADIY LAZAROV}, MARJATTA LYYRA, Department of Physics, Temple University, Philadelphia, PA 19122, USA; LI LI, Department of Modern Applied Physics, Tsinghua University, Beijing, 100084 CHINA; JOHN HUENNEKENS, Department of Physics, Lehigh University, 16 Memorial Dr. East, Bethlehem, PA 18015, USA.

The Na_{2} 4^{3}Π_{g} state^{a} has been studied by continuous wave Perturbation-Facilitated Optical-Optical Double Resonance (PFOODR) fluorescence excitation spectroscopy. The absolute vibrational numbering was determined by resolved fluorescence to the a^{3}Σ_{u}^{+} state. The hyperfine splitting has been resolved and analyzed. The hyperfine constant b_{F} is reported.

aY. Liu, J. Li, H. Chen, D. Chen, Li Li, and R. W. Field, to be submitted to J. Mol. Spectrosc.